OpenStudy (anonymous):

how to solve 6x^2-7x=12 by completing the square. Can someone show me how to do this??? I got many different answers and need to know which is correct.

OpenStudy (anonymous):

if you want to complete the square it will be somewhat of a pain because the leading coefficient is 6 we can still do it if you like

OpenStudy (anonymous):

it would probably be easiest to solve this using the quadratic formula, which completes the square automatically, but we can work through it step by step. your call

OpenStudy (anonymous):

which ever is easy I tried all ways and got 7+square root 337/12, 7-square root 337/12

OpenStudy (anonymous):

\[7+\sqrt{\frac{ 337 }{12 }},-7\sqrt{337/12}\]

OpenStudy (anonymous):

Ok, first get anything that includes the variable on one side